OpenStudy (anonymous):

the point (13,5) lies on the graph of 2log base b x. the value of b is

OpenStudy (anonymous):

@mertsj can u help me?

OpenStudy (anonymous):

\[2\log_{b} x\]

OpenStudy (anonymous):

can u help me with one more question after that? please?

OpenStudy (anonymous):

\[\log_{3b} t=M\] given this equation solve for b and give answer inexponential form?

OpenStudy (mertsj):

\[2\log_{b}x=y \] \[\log_{b}x^2=y \] \[b^y=x^2\] \[b^5=13^2\] \[(b^5)^{\frac{1}{5}}=(13^2)^{\frac{1}{5}}\] \[b=13^{\frac{2}{5}}\]

OpenStudy (mertsj):

I hope that's right. Do you spot any errors?

OpenStudy (anonymous):

i think its right. what about the second question i posted in th

OpenStudy (anonymous):

?

OpenStudy (anonymous):

given this equation solve for b and give answer in exponential form? look at same question above. can u help me @satellite73?

OpenStudy (anonymous):

(13,5) lies on the graph of \(2\log_b( x)\). the value of b is

OpenStudy (anonymous):

no the other question

OpenStudy (anonymous):

we know \[2\log_b(13)=5\] and therefore \[\log_b(13)=7.5\] thus \[b^{7.5}=13\]

OpenStudy (anonymous):

this one \[\log_{3b} t=M\]

OpenStudy (mertsj):